You have got a number line they start with zero, v, w, end with one

a. What is the whole in this case?

b. Into how many equal parts is this unit (whole) divided?

c. 2/3 of the whole is represented by which letter on the numberline?

d. You have a numberline that start with zero, x, y, z and end with one

Place 1/4, 2/4 and 3/4 on the numberline?



1
Expert's answer
2021-06-24T12:45:14-0400

Solution.



a) The whole in this case is a segment whose ends are 0 and 1.

b) There are 3 equal parts in this whole.

c) "W(\frac{2}{3})."

d) "X(\frac{1}{4}),Y(\frac{2}{4}),Z(\frac{3}{4})."